Henry's father is three times as old as henry. Six years ago, he was five times older. How old is henry

Answer by TimothyLamb(4379)

x = father age now
y = henry age now
---
x = 3y
x - 6 = 5(y - 6)
---
put the system of linear equations into standard form
---
x - 3y = 0
x - 6 = 5y - 30
---
x - 3y = 0
x - 5y = -24
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = father age now = 36
y = henry age now = 12
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php